# Noema formIng Cluster survEy (NICE): A Census of Star Formation and Cold Gas Properties in Massive protoclusters at 1.5 ¡ ¡ 4

### v2

@@ 3.2 Star formation main sequence in dense environments
The SFR and stellar mass of all UVJ-active cluster members (Williams et al., 2009) are shown in Fig. 4. Galaxies at different redshifts are scaled to = 2.5, following the SFMS normalization and redshift evolution from Schreiber et al. (2015). Despite significant scatter, the median values of these galaxies closely follow the main sequence in the field, in line with results from H emitters in protoclusters at 2.5 (e.g., Spiderweb, Pérez-Martínez et al., 2023).
@@
The CO-detected members, which are among the most massive with the highest SFRs, may indicate a greater abundance of gas fueling star formation in these galaxies. However, this trend is not observed in the CO detected H emitters in the Spiderweb protocluster (Pérez-Martínez et al., 2025), where SFRs are derived from the dust-corrected H fluxes, possibly due to the differences in protocluster selection criteria. In contrast, the less massive galaxies display slightly elevated star formation compared to field galaxies, a trend similar to that observed in H emitters within the young protocluster USS1558-003 at = 2.53 (Hayashi et al., 2012; Daikuhara et al., 2024).
@@
It is important to note that the cluster members were primarily selected based on photometric redshifts, which may include interlopers. This may affect the results shown in Fig. 4, especially at the low mass end, where the contamination is most significant. On the other hand, the high-mass end is less impacted, as it is dominated by CO-detected galaxies, which are spectroscopically confirmed members.
@@ 3.3 Build-up of massive galaxies in the dense cores
The field of view (FoV) of the NOEMA and ALMA observations, with full-width half maximums (FWHM) of 30″and 1′, respectively, is sufficiently large to encompass the area within the virial radius of the protoclusters. This indicates that we are witnessing the rapid build-up of massive galaxies in the cores of these protoclusters at cosmic noon.
@@ 3.4 Gas content
The gas masses of CO-detected NICE-COSMOS galaxies span the range = 1010.6 1011.7 M⊙. The gas-to-stellar mass ratio = / and gas depletion time = /SFR of NICE protocluster galaxies have median values of ,NICE = 1.32, and ,NICE = 0.78 Gyr. These values are consistent with those observed in other protocluster galaxies at similar redshifts (e.g., Tadaki et al., 2019; Pérez-Martínez et al., 2025). However, protoclusters at higher redshifts ( 4), such as SPT2349-56 (Hill et al., 2022) and LH-SBC3 (Zhou et al., 2024), tend to show lower and while massive galaxies in Distant Red Core at =4 (Long et al., 2020) have comparable gas content to those in protoclusters at 2.5. For a fair comparison, we recalculated the gas masses of galaxies in the literature, using the same method in Section 2.3 when CO(1-0) measurements were unavailable.
@@
To assess the environmental impact on the gas content, we normalized and to the corresponding field levels, accounting for their dependence on stellar mass, starburstiness and redshift. This analysis utilized parameterized relations compiled from the A3COSMOS project, based on a systematic mining of ALMA archival data for CO-observed galaxies in the COSMOS field (Liu et al., 2019). The results are presented in Fig. 6.
@@
The 3 detection limit of the NICE observations corresponds to a gas mass of 1011 M⊙ for galaxies at 2.5, meaning that our sample predominantly includes the most massive members and less massive galaxies with high starburstiness (Fig. 6c). The most massive ones ( ¿ 1010.8 M⊙), which are concentrated in the core regions, are nearly all detected via CO or dust continuum emission. These massive galaxies generally exhibit higher gas content (2) compared to their counterparts in the field (Fig. 6 a, b). Several of them display significantly elevated gas-to-stellar mass ratios and gas depletion times relative to field galaxies. However, their SFRs are not correspondingly enhanced, suggesting that they have abundant gas reservoirs to sustain their growth or their SFRs are prevented from increasing to starburst levels due to merger related processes or negative feedback.
@@
We also note that field galaxies (Kaasinen et al., 2019; Boogaard et al., 2020; Frias Castillo et al., 2023), whose molecular gas content is derived from CO(1–0) emission adopting obtained using the method outlined in Section 2.3, fall within the field-level range defined by the A3COSMOS project. This is consistent with findings from Calvi et al. (2023), who found a correlation between molecular gas masses of SMGs and the significance of the associated overdensity.
@@
Most galaxies with lower stellar masses are not detected in CO or dust continuum, except for the starbursts. However, these starburst galaxies exhibit gas-to-stellar mass ratios and gas depletion times comparable to those of field galaxies. It remains unclear whether the CO- or dust-non-detected cluster members have enhanced or deficient gas content.
@@
In contrast, three protoclusters traced by Ly emitters and/or red galaxies surrounding a radio galaxy (Tadaki et al., 2019) show the opposite trend, with a declining gas fraction and gas depletion time as stellar mass increases. This indicates a mass-dependent environmental effect, where gas accretion is accelerated in less massive galaxies while suppressed in more massive ones. These contrasting trends may reflect different evolutionary stages of the two protocluster populations.
@@
Interestingly, we find that most of the gas-rich galaxies belong to protoclusters at 2.5, whereas massive galaxies in more distant protoclusters exhibit a declining gas content, though still generally within the uncertainties of field galaxies (Fig. 6d,e) – approximately 2 times at = 2.5 and 1.2 times at = 4. Similarly, Calvi et al. (2023) identified a possible transitioning phase at 4, with ¡ 4 protoclusters more populated by dusty galaxies. We speculate that our observed trend may partly reflect a selection bias in constructing the field sample, which tends to favor the brightest sources at high redshifts, typically located in dense environments.
@@
We further checked the overall mean intensity of the radiation field, ¡¿, of the NICE-COSMOS protoclusters, as measured from the integrated FIR SED fitting in Sillassen et al. (2024). ¡¿ can be used to infer the metallicity weighted star formation efficiency (SFE 1/) and is proportional to dust temperature (Magdis et al., 2012, 2017). The evolution of ¡¿ with redshift for the eight protoclusters generally follow the trend observed in main sequence galaxies (Béthermin et al., 2015), with a few of them showing lower ¡¿. This is in line with the moderately longer gas depletion time observed in the individual member galaxies and may indicate a colder dust temperature in these protoclusters.
@@ 4 Summary and Discussion
In this paper, we analyzed the star formation and cold gas properties of eight protoclusters in the COSMOS field, selected through the NICE survey, which targets intensively star-forming protocluster cores at cosmic noon.
@@
We found that the total SFR per unit halo mass () in protoclusters at ¿ 2 increases rapidly with redshift, scaling as (Fig. 3). This could be due to either a higher number of massive star-forming galaxies in massive halos or enhanced star formation in cluster galaxies. Additionally, we find that the fraction of quiescent galaxies decreases relative to that observed in lower-redshift (proto)clusters, and that the streaming mass of the halos (i.e. the threshold mass delimiting the hot and cold in hot accretion regimes, see Daddi et al. 2022 and Dekel et al. 2009) is also a good predictor for the quiescent fraction, making the gas accretion mode a potential explanation for the quiescent fraction evolution across halos (Elias et al., in prep). We also caution that the higher values at higher redshifts could also be partly driven by lower halo masses.
@@
Nevertheless, this rise in star formation activity in protoclusters is not reflected in individual member galaxies. The star formation main sequence remains comparable to that observed in the field (Fig. 4), suggesting that the elevated in protoclusters is not driven by a higher fraction of starburst galaxies in dense environments. We note that the cluster members may be affected by contamination due to photometric redshift uncertainties (Sillassen et al., 2024). However, the massive end is dominated by CO-detected and thus spectroscopically confirmed members, making the main sequence at the high-mass end less susceptible to interlopers.
@@
Using observations from NOEMA and ALMA in the NICE survey, we investigated the gas content of cluster galaxies. The CO or dust continuum detected galaxies are predominantly massive cluster members ( ¿ 31010 M⊙), which preferentially reside in the core of the protoclusters ( 0.3, Fig. 5). This indicates an overabundance of massive star-forming galaxies in these protocluster cores, contributing to their elevated . This finding aligns with the simulations by Araya-Araya et al. (2024), which identify an excess of massive star-forming galaxies undergoing a submillimeter-bright phase in protocluster cores, suggesting a ”downsizing” formation scenario for massive galaxies.
@@
In terms of stellar-to-gas mass ratio and gas depletion time, the median values are consistent with field levels. However, the most massive galaxies ( ¿ 81010 M⊙) are notably gas-richer than their field counterparts (Fig. 6). This suggests efficient cold gas accretion in the central regions of these protoclusters, supporting the growth of these galaxies to higher masses while remaining in the main sequence mode, without triggering starburst phases. For the less massive galaxies, CO or dust continuum detections are limited to the most starbursting ones, which follow the scaling relation ((, , ), (, , ), Liu et al., 2019) observed in the field galaxies. The moderately higher is consistent with the overall mean intensity of the radiation field in the protoclusters, indicating a generally cold dust temperature in these dense environments. However, we caution that the NICE survey is biased toward protoclusters in active star-forming phases. The influence of such environments on member galaxies may differ from that observed in protoclusters selected through other methods.
@@ Acknowledgements.
We thank the anonymous referee for valuable comments. This work was supported by National Natural Science Foundation of China (Project Nos. 13001103, 12173017 and Key Project No.12141301), National Key R&D Program of China (grant no. 2023YFA1605600), Scientific Research Innovation Capability Support Project for Young Faculty (Project No. ZYGXQNJSKYCXNLZCXM-P3), and the China Manned Space Program with grant no. CMS-CSST-2025-A04. L.Z. and Y.S. acknowledges the support from the National Key R&D Program of China No. 2022YFF0503401 and 2018YFA0404502, the National Natural Science Foundation of China (NSFC grants 12141301, 12121003, 11825302). Y.S. thanks the support by the New Cornerstone Science Foundation through the XPLORER PRIZE. Y.J.W. acknowledges support by National Natural Science Foundation of China (Project No. 12403019) and Jiangsu Natural Science Foundation (Project No. BK20241188). GEM acknowledges the Villum Fonden research grant 13160 “Gas to stars, stars to dust: tracing star formation across cosmic time,” grant 37440, “The Hidden Cosmos,” and the Cosmic Dawn Center of Excellence funded by the Danish National Research Foundation under the grant No. 140. CGG acknowledges support from CNES. ZJ acknowledges funding from JWST/NIRCam contract to the University of Arizona NAS5-02015. ID acknowledges funding by the INAF Minigrant ”Harnessing the power of VLBA towards a census of AGN and star formation at high redshift” and by the European Union – NextGenerationEU, RRF M4C2 1.1, Project 2022JZJBHM: ”AGN-sCAN: zooming-in on the AGN-galaxy connection since the cosmic noon” - CUP C53D23001120006. CdE acknowledges funding from the MCIN/AEI (Spain) and the “NextGenerationEU”/PRTR (European Union) through the Juan de la Cierva-Formación program (FJC2021-047307-I). SJ is supported by the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101060888. This research used APLpy, an open-source plotting package for Python (Robitaille & Bressert, 2012; Robitaille, 2019). This work is based on observations carried out under project number M21AA with the IRAM NOEMA Interferometer. IRAM is supported by INSU/CNRS (France), MPG (Germany) and IGN (Spain). This research uses data obtained through the Telescope Access Program (TAP), which is funded by the National Astronomical Observatories, Chinese Academy of Sciences, and the Special Fund for Astronomy from the Ministry of Finance. This work is based on observations obtained with WIRCam, a joint project of CFHT, Taiwan, Korea, Canada, France, at the Canada-France-Hawaii Telescope (CFHT) which is operated by the National Research Council (NRC) of Canada, the Institute National des Sciences de l’Univers of the Centre National de la Recherche Scientifique of France, and the University of Hawaii.References
@@
- Pérez-Martínez et al. (2025) Pérez-Martínez, J. M., Dannerbauer, H., Emonts, B. H. C., et al. 2025, A&A, 696, A236
- Pérez-Martínez et al. (2023) Pérez-Martínez, J. M., Dannerbauer, H., Kodama, T., et al. 2023, MNRAS, 518, 1707

### v3

@@ 3.2 Star-forming main sequence in dense environments
The SFR and stellar mass of all UVJ-active cluster members (Williams et al., 2009) are shown in Fig. 4. Galaxies at different redshifts are scaled to = 2.5, following the SFMS normalization and redshift evolution from Schreiber et al. (2015). Despite significant scatter, the median values of these galaxies closely follow the main sequence in the field, in line with the results from H emitters in protoclusters at 2.5 (e.g., Spiderweb, Pérez-Mart´ınez et al., 2023).
@@
The CO-detected members, which are among the most massive with the highest SFRs, may indicate a greater abundance of gas fueling star formation in these galaxies. However, this trend is not observed in the CO-detected H emitters in the Spiderweb protocluster (Pérez-Mart´ınez et al., 2025), where SFRs are derived from the dust-corrected H fluxes, possibly due to the differences in protocluster selection criteria. In contrast, the less massive galaxies display slightly elevated star formation compared to field galaxies, a trend similar to that observed in H emitters within the young protocluster USS1558-003 at = 2.53 (Hayashi et al., 2012; Daikuhara et al., 2024).
@@
It is important to note that the cluster members were primarily selected based on photometric redshifts, which may include interlopers. This may affect the results shown in Fig. 4, especially at the low-mass end, where the contamination is most significant. On the other hand, the high-mass end is less impacted as it is dominated by CO-detected galaxies that are spectroscopically confirmed members.
@@ 3.3 Build-up of massive galaxies in the dense cores
The field of view (FoV) of the NOEMA and ALMA observations, with a full width at half maximum (FWHM) of 30″and 1′, respectively, is sufficiently large to encompass the area within the virial radius of the protoclusters. This indicates that we are witnessing the rapid build-up of massive galaxies in the cores of these protoclusters at cosmic noon.
@@ 3.4 Gas content
The gas masses of CO-detected NICE-COSMOS galaxies span the range = 1010.6 1011.7 M⊙. The gas-to-stellar mass ratio = / and gas depletion time = /SFR of NICE protocluster galaxies have median values of ,NICE = 1.32and ,NICE = 0.78 Gyr. These values are consistent with those observed in other protocluster galaxies at similar redshifts (e.g., Tadaki et al., 2019; Pérez-Mart´ınez et al., 2025). However, protoclusters at higher redshifts ( 4), such as SPT2349-56 (Hill et al., 2022) and LH-SBC3 (Zhou et al., 2024), tend to show lower and , while massive galaxies in Distant Red Core at =4 (Long et al., 2020) have comparable gas content to those in protoclusters at 2.5. For a fair comparison, we recalculated the gas masses of galaxies in the literature using the same method described in Section 2.3 when CO(1-0) measurements were unavailable.
@@
To assess the environmental impact on the gas content, we normalized and to the corresponding field levels, accounting for their dependence on stellar mass, starburstiness, and redshift. This analysis utilized parameterized relations compiled from the A3COSMOS project, based on a systematic mining of ALMA archival data for CO-observed galaxies in the COSMOS field (Liu et al., 2019). The results are presented in Fig. 6.
@@
The 3 detection limit of the NICE observations corresponds to a gas mass of 1011 M⊙ for galaxies at 2.5, meaning that our sample predominantly includes the most massive members and the less massive galaxies with high starburstiness (Fig. 6c). The most massive ones ( ¿ 1010.8 M⊙), which are concentrated in the core regions, are nearly all detected via CO or dust continuum emission. These massive galaxies generally exhibit higher gas content (2) compared to their counterparts in the field (Fig. 6 a, b). Several of them display significantly elevated gas-to-stellar mass ratios and gas depletion times relative to field galaxies. However, their SFRs are not correspondingly enhanced, suggesting that they have abundant gas reservoirs to sustain their growth or their SFRs are prevented from increasing to starburst levels due to merger-related processes or negative feedback.
@@
We also note that the field galaxies (Kaasinen et al., 2019; Boogaard et al., 2020; Frias Castillo et al., 2023), whose molecular gas content is derived from CO(1–0) emission adopting obtained using the method outlined in Section 2.3, fall within the field-level range defined by the A3COSMOS project. This is consistent with the findings from Calvi et al. (2023), who found a correlation between molecular gas masses of SMGs and the significance of the associated overdensity.
@@
Most galaxies with lower stellar masses are not detected in CO or dust continuum, except for the starbursts. However, these starburst galaxies exhibit gas-to-stellar mass ratios and gas depletion times comparable to those of field galaxies. It remains unclear whether the CO- or dust-nondetected cluster members have enhanced or deficient gas content.
@@
In contrast, three protoclusters traced by Ly emitters and/or red galaxies surrounding a radio galaxy (Tadaki et al., 2019) show the opposite trend, with a declining gas fraction and gas depletion time as stellar mass increases. This indicates a mass-dependent environmental effect, where gas accretion is accelerated in less massive galaxies, while suppressed in more massive ones. These contrasting trends may reflect different evolutionary stages of the two protocluster populations.
@@
Interestingly, we find that most of the gas-rich galaxies belong to protoclusters at 2.5, whereas massive galaxies in more distant protoclusters exhibit a declining gas content, though still generally within the uncertainties of field galaxies (Fig. 6d,e), approximately 2 times at = 2.5 and 1.2 times at = 4. Similarly, Calvi et al. (2023) identified a possible transitioning phase at 4, with ¡ 4 protoclusters more populated by dusty galaxies. We speculate that our observed trend may partly reflect a selection bias in constructing the field sample, which tends to favor the brightest sources at high redshifts, typically located in dense environments.
@@
We further checked the overall mean intensity of the radiation field, ¡¿, of the NICE-COSMOS protoclusters, as measured from the integrated FIR SED fitting in Sillassen et al. (2024). ¡¿ can be used to infer the metallicity weighted star formation efficiency (SFE 1/) and is proportional to dust temperature (Magdis et al., 2012, 2017). The evolution of ¡¿ with redshift for the eight protoclusters generally follows the trend observed in main sequence galaxies (Béthermin et al., 2015); a few of them show lower ¡¿. This is in line with the moderately longer gas depletion time observed in the individual member galaxies and may indicate a colder dust temperature in these protoclusters.
@@ 4 Summary and discussion
In this paper we analyzed the star formation and cold gas properties of eight protoclusters in the COSMOS field, selected through the NICE survey, which targets intensively star-forming protocluster cores at cosmic noon.
@@
We find that the total SFR per unit halo mass () in protoclusters at ¿ 2 increases rapidly with redshift, scaling as (Fig. 3). This could be due to either a higher number of massive star-forming galaxies in massive halos or enhanced star formation in cluster galaxies. Additionally, we find that the fraction of quiescent galaxies decreases relative to that observed in lower-redshift (proto)clusters, and that the streaming mass of the halos (i.e., the threshold mass delimiting the hot and cold in hot accretion regimes; see Daddi et al. 2022 and Dekel et al. 2009) is also a good predictor for the quiescent fraction, making the gas accretion mode a potential explanation for the quiescent fraction evolution across halos (Elias et al. in prep.). We also caution that the higher values at higher redshifts could also be partly driven by lower halo masses.
@@
Nevertheless, this rise in star formation activity in protoclusters is not reflected in individual member galaxies. The SFMS remains comparable to that observed in the field (Fig. 4), suggesting that the elevated in protoclusters is not driven by a higher fraction of starburst galaxies in dense environments. We note that the cluster members may be affected by contamination due to photometric redshift uncertainties (Sillassen et al., 2024). However, the massive end is dominated by CO-detected and thus spectroscopically confirmed members, making the main sequence at the high-mass end less susceptible to interlopers.
@@
Using observations from NOEMA and ALMA in the NICE survey, we investigated the gas content of cluster galaxies. The CO or dust continuum-detected galaxies are predominantly massive cluster members ( ¿ 31010 M⊙), which preferentially reside in the core of the protoclusters ( 0.3, Fig. 5). This indicates an overabundance of massive star-forming galaxies in these protocluster cores, which contributes to their elevated . This finding aligns with the simulations by Araya-Araya et al. (2024), which identify an excess of massive star-forming galaxies undergoing a submillimeter-bright phase in protocluster cores, suggesting a downsizing formation scenario for massive galaxies.
@@
In terms of stellar-to-gas mass ratio and gas depletion time, the median values are consistent with field levels. However, the most massive galaxies ( ¿ 81010 M⊙) are notably more gas-rich than their field counterparts (Fig. 6). This suggests efficient cold gas accretion in the central regions of these protoclusters, supporting the growth of these galaxies to higher masses while remaining in the main sequence mode, without triggering starburst phases. For the less massive galaxies, CO or dust continuum detections are limited to the most starbursting ones, which follow the scaling relation ((, , ), (, , ), Liu et al., 2019) observed in the field galaxies. The moderately higher is consistent with the overall mean intensity of the radiation field in the protoclusters, indicating a generally cold dust temperature in these dense environments. However, we caution that the NICE survey is biased toward protoclusters in active star-forming phases. The influence of such environments on member galaxies may differ from that observed in protoclusters selected through other methods.
@@ Acknowledgements.
We thank the anonymous referee for valuable comments. This work was supported by National Natural Science Foundation of China (Project Nos. 13001103, 12173017 and Key Project No.12141301), National Key R&D Program of China (grant no. 2023YFA1605600), Scientific Research Innovation Capability Support Project for Young Faculty (Project No. ZYGXQNJSKYCXNLZCXM-P3), and the China Manned Space Program with grant no. CMS-CSST-2025-A04. L.Z. and Y.S. acknowledges the support from the National Key R&D Program of China No. 2022YFF0503401 and 2018YFA0404502, the National Natural Science Foundation of China (NSFC grants 12141301, 12121003, 11825302). Y.S. thanks the support by the New Cornerstone Science Foundation through the XPLORER PRIZE. Y.J.W. acknowledges support by National Natural Science Foundation of China (Project No. 12403019) and Jiangsu Natural Science Foundation (Project No. BK20241188). GEM acknowledges the Villum Fonden research grant 13160 “Gas to stars, stars to dust: tracing star formation across cosmic time,” grant 37440, “The Hidden Cosmos,” and the Cosmic Dawn Center of Excellence funded by the Danish National Research Foundation under the grant No. 140. CGG acknowledges support from CNES. ZJ acknowledges funding from JWST/NIRCam contract to the University of Arizona NAS5-02015. ID acknowledges funding by the INAF Minigrant ”Harnessing the power of VLBA toward a census of AGN and star formation at high redshift” and by the European Union – NextGenerationEU, RRF M4C2 1.1, Project 2022JZJBHM: ”AGN-sCAN: zooming-in on the AGN-galaxy connection since the cosmic noon” - CUP C53D23001120006. CdE acknowledges funding from the MCIN/AEI (Spain) and the “NextGenerationEU”/PRTR (European Union) through the Juan de la Cierva-Formación program (FJC2021-047307-I). SJ is supported by the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101060888. This research used APLpy, an open-source plotting package for Python (Robitaille & Bressert, 2012; Robitaille, 2019). This work is based on observations carried out under project number M21AA with the IRAM NOEMA Interferometer. IRAM is supported by INSU/CNRS (France), MPG (Germany) and IGN (Spain). This research uses data obtained through the Telescope Access Program (TAP), which is funded by the National Astronomical Observatories, Chinese Academy of Sciences, and the Special Fund for Astronomy from the Ministry of Finance. This work is based on observations obtained with WIRCam, a joint project of CFHT, Taiwan, Korea, Canada, France, at the Canada-France-Hawaii Telescope (CFHT) which is operated by the National Research Council (NRC) of Canada, the Institute National des Sciences de l’Univers of the Centre National de la Recherche Scientifique of France, and the University of Hawaii.References
@@
- Pérez-Mart´ınez et al. (2025) Pérez-Martínez, J. M., Dannerbauer, H., Emonts, B. H. C., et al. 2025, A&A, 696, A236
- Pérez-Mart´ınez et al. (2023) Pérez-Martínez, J. M., Dannerbauer, H., Kodama, T., et al. 2023, MNRAS, 518, 1707